# Ranking Unraveled: Recipes for LLM Rankings in Head-to-Head AI Combat

### v1

@@ Abstract
Deciding which large language model (LLM) to use is a complex challenge. Pairwise ranking has emerged as a new method for evaluating human preferences for LLMs. This approach entails humans evaluating pairs of model outputs based on a predefined criterion. By collecting these comparisons, a ranking can be constructed using methods such as Elo. However, applying these algorithms as constructed in the context of LLM evaluation introduces several challenges. In this paper, we explore the effectiveness of ranking systems for head-to-head comparisons of LLMs. We formally define a set of fundamental principles for effective ranking and conduct a series of extensive evaluations on the robustness of several ranking algorithms in the context of LLMs. Our analysis uncovers key insights into the factors that affect ranking accuracy and efficiency, offering guidelines for selecting the most appropriate methods based on specific evaluation contexts and resource constraints.
@@ 1 Introduction
Figure 1: Ranking algorithms can produce different ranks from the same human evaluation data making it difficult to determine which algorithm is appropriate for various use-case.
The adoption of Large Language Models (LLMs) has surged in recent years, revolutionizing natural language processing (NLP) tasks across various domains. A critical question arises when practitioners use these models “Which LLM is most suitable for a particular task or use case?" Given the plethora of LLMs available today, each with their own unique strengths and weaknesses, this decision-making process has become increasingly complex Jiang et al. (2023).
Historically, NLP benchmarks such as GLUE Wang et al. (2019), SuperGLUE Wang et al. (2020), and more recently LLM-centered benchmarks such as LM-Eval Gao et al. (2023) have aimed to answer this question by evaluating these models on a series of challenging tasks that assess their natural language understanding (NLU) capabilities. However, a major limitation of these classical benchmarks is that they often fail to capture and evaluate more nuanced, diverse, and complex cases where establishing a definitive ground truth is challenging and, in many cases, unattainable Chiang et al. (2024). Consider the task of rating competing model responses to a given prompt. These standard benchmarks often overlook the intricacies of assessing language generation, where nuances and qualitative factors play an important role. This problem is particularly evident in tasks such as open-ended text generation, conversational dialogue, and creative writing, where alignment with human judgment is paramount. Although a given model may score highly on these classical benchmarks, their output often does not align with human preference. Zheng et al. (2024); Irugalbandara et al. (2023).
To better align with human preference, open benchmarking platforms such as Chatbot Arena Chiang et al. (2024) evaluate models through "head-to-head combat," where models are judged and ranked individually inspired by the competitive dynamics of sports leagues. This approach assesses the relative strengths of LLMs through confrontations, enabling a more accurate comparison between them. This pairwise comparison between models naturally aligns with the Elo rating system Elo (1978), originally developed to classify chess players according to their relative strengths. When evaluating the performance of the LLMs in pairwise competition, Elo and ranking methods alike have gained popularity as structured methods to incorporate human feedback and preferences to effectively rank models Dettmers et al. (2024); Wu and Aji (2023); Chiang et al. (2023). However, despite its recent prominence in LLM evaluation, Elo can produce inconsistent and sometimes unreliable results Boubdir et al. (2023). Elo was designed for domains with stable and well-defined results, such as chess. LLMs, in contrast, are different as their performance can be multifaceted and context-dependent, making Elo less than ideal for capturing the full scope of their capabilities. This added complexity raises several important considerations when determining the ranking system for model selection. Choosing the best model can have significant implications for the user experience of the application that incorporates it.
This paper explores the critical issues around ranking LLMs in head-to-head combat. We first outline a set of fundamental properties that ranking systems should adhere to when evaluating LLMs. Second, we evaluate the performance of four widely used ranking algorithms (Elo, Bradley-Terry, Glicko, and Markov Chain) in two different evaluation scenarios (Arena Style vs. Controlled Style), analyzing their ability to maintain transitivity, ensure prediction accuracy, and remain stable across hyperparameter variations. Lastly, we conduct an in-depth analysis of these results, presenting a series of insights into the factors influencing the performance of different ranking methodologies and propose a set of guidelines for selecting the most appropriate approach based on the specific characteristics of the evaluation task and available resources. To the best of our knowledge, this paper represents the first work that systematically addresses the complexities and intricacies of ranking LLMs across various methodologies. We release all code, data, and models to facilitate reproducibility and further research in this area [^1] http://github.com/daynauth/elo_paper.
@@ 2 Background
In this section, we formally define the task of pairwise ranking and introduce the suite of ranking algorithms considered in our work.
@@ 2.1 Ranking
If $F$ follows a logistic distribution, then the system is known as the Bradley-Terry model Bradley and Terry (1952) and
$p_{ij}=\frac{1}{1+e^{-(\theta_{i}-\theta_{j})}}$ (2)
@@ 2.3 Bradley-Terry Model
where $y_{ij}$ is the number of times that model $i$ beats $j$ and $p_{ij}$ is the probability of $i$ beating $j$. This process is optimized by the maximum likelihood of $\log\mathcal{L}$, i.e, solving $\log\mathcal{L(\theta)}=\sum_{i>j}y_{ij}\log(p_{ij})+(1-y_{ij})\log(1-p_{ij})$.
@@ 2.4 Glicko Rating system
The Glicko rating system Glickman (1999) was introduced in 1995 as an improvement to Elo. Glicko extends Elo by introducing a secondary parameter $(\sigma)$ known as the rating deviation. This parameter measures the reliability of a player’s rating score, acting as a confidence interval for the player’s rank, decreasing based on the number of games played, regardless of wins or losses. This factor is important given that two players can potentially have the same score in a tournament despite a different number of previous matches. This situation is prevalent in LLM evaluations, as new models are frequently introduced and the number of games played between them vary significantly.

### v2

@@ Abstract
Evaluating large language models (LLMs) is a complex task. Pairwise ranking, where humans compare LLM outputs based on predefined criteria, has become a leading approach. By aggregating these comparisons through algorithms such as Elo, rankings across multiple LLMs can be derived. However, applying ranking algorithms in LLM evaluation presents several challenges. Traditional systems like Elo, designed initially for structured competitions such as chess, often produce inconsistent and unstable rankings due to the dynamic and context-dependent nature of LLM performance. Despite the increasing reliance on these methods, a systematic study of ranking algorithms for LLM evaluation remains lacking. This paper examines the effectiveness of various ranking systems for head-to-head LLM comparisons. We define key principles for robust ranking, conduct extensive evaluations of different ranking algorithms, and analyze their stability, accuracy, and sensitivity to real-world conditions. Our findings offer insights into the limitations of existing approaches and provide guidelines for selecting the most appropriate ranking method based on evaluation objectives and resource constraints.
@@ 1 Introduction
Figure 1: Different ranking algorithms can produce different rankings with the same human evaluation data, making it difficult to determine which algorithm is appropriate for various use cases.
Rapid progress in large language models (LLMs) has highlighted the critical need for LLM evaluation methodologies, especially for complex generation tasks Guo et al. (2023). Traditionally, NLP evaluation has relied on benchmarks such as GLUE Wang et al. (2019), SuperGLUE Wang et al. (2020), and more recent LLM-focused benchmarks like LM-Eval Gao et al. (2023). However, a significant limitation of these classical benchmarks is their reliance on established ground truth, making them unsuitable for evaluating more nuanced, diverse, and complex cases where such ground truth is difficult to define or entirely unattainable Chiang et al. (2024). This problem is particularly evident in complex tasks such as open-ended text generation, conversational dialogue, and creative writing, where human judgment is paramount. Previous work has shown that these benchmark evaluation results do not correlate well with human LLM evaluations Zheng et al. (2024); Irugalbandara et al. (2023).
This shortcoming has led the NLP community to adopt pairwise human evaluation methods Dubois et al. (2024); Zheng et al. (2024); Chiang et al. (2024). A widely recognized example is Chatbot Arena Chiang et al. (2024), which directly compares models through "head-to-head combat," where users vote between responses from two competing models. This method aligns with the Elo rating system Elo (1978), initially developed to classify chess players according to their historical competition results and relative strengths. Elo and other ranking methods have gained popularity as structured methods incorporating human feedback and preferences to effectively rank models Dettmers et al. (2024); Wu and Aji (2023); Chiang et al. (2023).
However, as our experiments later demonstrated, Elo and other established ranking algorithms can often produce inconsistent and sometimes unreliable results (Figure 1), generating conflicting ranks. Thus, it is critical to understand how to select the ranking algorithm appropriately. There has been a lack of comprehensive studies across multiple ranking algorithms applied to LLM evaluations.
  • First, a general framework for such a study has been lacking. It remains unclear which key metrics should be used to evaluate ranking algorithms and the factors that can affect the accuracy of the ranking algorithm.
  • There has been no comprehensive study comparing multiple ranking algorithms or offering guidelines for selecting the most suitable in the context of LLM evaluations. Although prior work such as Boubdir et al. (2023) has explored some challenges associated with using Elo for LLM ranking, its focus is limited to Elo. In addition, our experiments show that recommendations from previous work are not always effective in practice.
Choosing the best model can significantly impact the user experience of the application that incorporates it. This paper explores the critical issues surrounding ranking LLMs in head-to-head human evaluations. Our contributions are as follows.
  • We first systemically outline a set of fundamental properties that ranking systems should adhere to when evaluating LLMs.
  • Second, we evaluate the performance of four widely used ranking algorithms (Elo, Bradley-Terry, Glicko, and Markov Chain) in two different evaluation scenarios (Arena Style vs. Controlled Style), analyzing their ability to maintain transitivity, ensure prediction accuracy, and remain stable across hyperparameter variations.
  • Third, we conduct an in-depth analysis of these results, highlighting flaws in the current application of these popular algorithms. Despiteits wide use for LLM ranking, Elo remains unstable even after extensive tuning is applied.
  • Lastly, we propose systematic guidelines for selecting the most appropriate approach based on the specific characteristics of the evaluation task and the available resources.
To our knowledge, this paper represents the first work that systematically addresses the complexities and intricacies of ranking LLMs across various methodologies. We release all code, data, and models to facilitate reproducibility and further research in this area.
@@ 2 Background
In this section, we formally define the task of pairwise ranking and briefly describe the suite of algorithms often used for aggregating ranking from pair-wise comparison results. The first three algorithms—Elo Elo (1978), Bradley-Terry Bradley and Terry (1952), and Glicko Glickman (1999)—are well-established and widely used in contexts such as sports competitions, while the Markov chain, based on a random walker model, is simpler in design and has demonstrated strong performance in previous studies Callaghan et al. (2003, 2004).
@@ 2.3 Bradley-Terry Model
where $y_{ij}$ is the number of times that model $i$ beats $j$ and $p_{ij}$ is the probability of $i$ beating $j$.
@@ 2.4 Glicko Rating system
The Glicko rating system Glickman (1999) was introduced in 1995 as an improvement to Elo. Glicko extends Elo by introducing a secondary parameter $(\sigma)$ known as the rating deviation. This parameter measures the reliability of a player’s rating score, acting as a confidence interval for the player’s rank, decreasing based on the number of games played, regardless of wins or losses. This factor is important given that two players can potentially have the same score in a tournament despite a different number of previous matches. This situation is prevalent in LLM evaluations, as new models are frequently introduced and the number of games played between them vary significantly. In Appendix A.1 we explore the details of the algorithm.
2.5 Markov Chain
The Markov Chain Model is a nonparametric ranking algorithm adapted from Callaghan et al. (2003, 2004) and adapted into a Markov Chain Model by Kvam and Sokol (2006) (see Appendix A.2 for details). The Markov chain employs a series of random walkers traversing a graph, with each node representing a player/model in the tournament connected by an edge denoting a match between two nodes. Each walker chooses an initial node $i$, staying in the node or moving to an adjacent node $j$ according to a single condition $p$. In a random match between $i$ and $j$, the walker moves to the winning node with a probability of $p$ for $p>0.5$, a hyperparameter chosen in advance. Every time a walker moves or stays on a node, it represents a vote for that node, and the sum of all votes from every walker represents the rank of that player/model. In practice, the optimal value for $p$ can vary depending on the tournament. In our experiments, we use $p=0.8$ as the default value similar to Callaghan et al. (2003), which achieved stable results with $p$ ranging from $p=0.75$ to $p=0.95$.
3 Desirable Properties of Ranking System
The goal of a ranking system is to provide a structured approach to compare and order items based on their performance or quality. In the context of LLMs, an effective ranking system allows researchers, practitioners, and end-users to identify the most capable models for their specific needs. . By simplifying decision-making and optimizing resource allocation, such a system highlights models that are likely to deliver superior performance in the real-world applications. However, implementing a robust ranking system requires addressing several key properties to ensure its validity and reliability. In this section, we identify and discuss three essential properties that a ranking system for LLMs should adhere to: 1) Transitivity, 2) Prediction Accuracy, and 3) Sensitivity to Hyperparameters and Battle Conditions.
3.1 Transitivity
Transitivity is a fundamental property for a ranking system to maintain consistency.
Definition 3.1.
For any three models $i,j,k\in M$; their match-ups are transitive if $p_{ij}>0.5,p_{jk}>0.5$ and $p_{ik}>0.5$. If $i,j,k$ are transitive, we say that a ranking system preserves transitivity if it ranks these models in order $i,j,k$. Transitivity ensures that the ranking is coherent and interpretable.
Human evaluation data contains transitive and intransitive relationships, whereas ranking algorithms always produce perfectly transitive rankings. Transitive and intransitive relationships are mutually exclusive; however, intransitive relationships are inevitably lost, since ranking enforces transitivity. One of the key strengths of a ranking algorithm lies in its ability to preserve as much of the original transitive structure from human data as possible. Failure to do so can lead to paradoxical rankings, where the ordering of models becomes inconsistent, ultimately diminishing trust in the rating system.
3.2 Prediction Accuracy
Like transitivity, prediction accuracy is a critical metric that assesses how well the ranking system forecasts the results in head-to-head matches. It measures the probability that the ranking system correctly predicts the winner between any two LLMs. High prediction accuracy indicates that the ranking system is reliable and can be trusted to make informed decisions about the superiority of the model for a downstream task. For combat-style evaluation in LLM, where assessment is based on subjective human judgment, prediction accuracy acts as an indicator of the ranking system’s ability to align with human preferences. This alignment is crucial for tasks like open-ended text generation or conversational agents, where the model’s performance is best judged by human preference. Thus, the ranking system should reflect statistical superiority and resonate with qualitative human evaluations to truly enhance the utility and adoption of LLM.
3.3 Sensitivity to Hyperparameters
4 Evaluating Ranking Algorithms
Algorithm Arena (%) SLAM (%)
Elo 68.24 52.5
Markov 51.38 51.67
Glicko 56.54 53.33
Bradley-Terry 77.29 56.67
4.1 Datasets
To evaluate these ranking systems, we use two diverse human evaluation datasets representative of two evaluation styles:
These two diverse datasets enable us to assess ranking algorithms under varying conditions, offering a comprehensive view of their behavior. To evaluate dataset consistency, we apply Kendall’s test Kendall and Smith (1940) (see A.6 for details), which confirms that any inconsistencies in the datasets are statistically insignificant.
4.2 Preserving Transitivity
Using the Chatbot Arena and SLAM datasets, we evaluated the ability of each ranking algorithm to preserve transitivity (see 3.1). We evaluated the transitive performance of each ranking algorithm by first calculating the number of triples within the data set. Three models $i,j,k$ are part of a triple if $i$ wins the majority of comparisons with $j$, and $j$ wins the majority of comparisons with $k$, and transitivity is considered preserved if this order is maintained in the final ranking. Table 1 shows the performance of each ranking algorithm under the arena and controlled style evaluation scenarios. We observe that the Bradley-Terry model is the most performant, preserving transitivity in $77.29\%$ and $56.67\%$ of the triples, respectively, compared to its closest competitor, Elo, at $68.24\%$ and $52.5\%$.
Bradley-Terry consistently outperforms other algorithms in preserving transitivity, achieving the highest rankings across both datasets.
Algorithm Arena (F1) SLAM (F1)
Elo 0.90 0.87
Markov 0.77 0.88
Glicko 0.88 0.88
Bradley-Terry 0.82 0.87
Table 2: Ranking algorithm performance in prediction accuracy (F1 score) on Chatbot Arena & SLAM Dataset.
4.3 Prediction Accuracy
Performance on the Arena Dataset
Due to its uneven distribution of model match-ups, the Arena data set presents an interesting challenge for ranking algorithms to predict future outcomes. When matches between models are unevenly distributed, some algorithms struggle with prediction accuracy. Our results show the Elo model to be the most accurate, achieving an F1 score of $0.90$ compared to the F1 score of $0.88$, $0.82$, and $0.77$ for the Glicko, Bradley-Terry and Markov Chain models, respectively (Table 2).
The impact of the match-up distribution on the prediction accuracy is shown in the performance of the Markov chain which is affected by the sparsity of certain model match-ups, leading to lower prediction accuracy, as shown by the poor performance on models such as deepseek-llm-67b-chat and dolphin-2.2.1-mistral-7b in figure 4 of the appendix.
Another example of the impact of matchup distribution is shown in the performance of the Bradley-Terry model in the appendix figure 4. We observe that the Bradley-Terry system suffers when dealing with "powerful models," where there is a strong imbalance between a model’s wins and losses. For example, in the arena dataset, gpt-4-turbo had a win/loss ratio of 12288/3979. This skew in model wins/loss results in overestimation of the model’s strength as shown by the low F1 score of $0.82$ achieved by gpt-4-turbo using Bradley-Terry compared to an F1 score of $0.95$, $0.96$ and $0.92$ from Elo, Glicko, and Markov Chain, respectively. This is known as the ‘rare events’ problem in logistic regression King and Zeng (2001). One method of countering this is to use a weighted logistic regressionChiang et al. (2024). However, we observe that this has a negligible impact on the final result. Furthermore, the bootstrap confidence interval used by Chatbot Arena shows a narrow range of [-2.2, 2.3], suggesting a high confidence in the model ranking. However, this interval does not guarantee how well the model will perform against unseen opponents.
Prediction accuracy on the SLAM dataset
All algorithms perform similarly on the SLAM dataset, but on Arena, Elo leads in accuracy, followed by Glicko. Bradley-Terry, however, is negatively affected by the uneven distribution of the Arena dataset.
4.4 Hyperparameter Sensitivity
To measure the sensitivity of the hyperparameters, we calculate the F1 score of each ranking system as described in Section 4.3, varying each hyperparameter over a range of 100 values. Figure 2 showcases the range of results produced by Elo, Markov and Glicko across the two evaluation scenarios.
Method Elo Markov Glicko Bradley-Terry Win-Rate
Elo Rating 1.00/1.00 0.74/0.89 0.86/0.93 0.94/0.95 0.93/0.91
Markov - 1.00/1.00 0.76/0.99 0.81/0.97 0.78/0.98
Glicko - - 1.00/1.00 0.81/0.99 0.89/0.99
Bradley-Terry - - - 1.00/1.00 0.91/0.98
Win-Rate - - - - 1.00/1.00
Figure 2: Distribution of F1 scores for the SLAM (left) and ARENA (right) datasets, showing the performance of Elo, Markov, and Glicko algorithms across a subset of models with 100 different hyperparameter settings. The results highlight the volatility of Elo, demonstrating its high sensitivity to hyperparameter changes compared to the more stable performance of Markov and Glicko.
Elo k-factor
The $k$ factor in Elo dictates the extent to which the score of a player is influenced by the outcome of an individual match. For example, with large $k$ values, highly ranked players can incur a significant penalty when losing to a player with a lower ranking, while defeating such a player only results in a slight increase in score. The evaluation of the performance of the model (depicted in Figure 2) at varying levels of hyperparameter $k$ shows marked sensitivity, which emphasizes the need for meticulous tuning. Elo shows the largest variation across both datasets, and is less stable, particularly for the smaller SLAM dataset. Through our tuning of this parameter, we determined that the ideal $k$ factor for Elo is less than the standard 32 commonly used in previous research Dettmers et al. (2024); Wu and Aji (2023); Chiang et al. (2023). This reduced convergence rate can be attributed to the complexities of human preferences as opposed to the typical assessment of real-world games, which rely on the differing abilities of opposing teams.
Markov Chain: $p$ Score
The $p$ score in the Markov Chain determines the extent to which the overall ranking is influenced by each model’s win rate, with values closer to $1$ indicating greater influence compared to those closer to $0.5$. We observe that Markov shows stable performance on the controlled SLAM datasets, performing as well as Glicko (Figure 2), however, the algorithm struggles on the large ARENA dataset due to its over reliance on win-rates for calculating ranks. When optimizing the value of this parameter, we find that the optimal $p$ value for the Markov chain on the SLAM dataset is larger than the Arena dataset, indicating that the model’s strength is greatly influenced by its overall win rate.
Glicko: Rating deviation
For Glicko, the ranking deviation $(\sigma)$ is a measure of consistency or variability in a player’s performance over time and determines how flexible a player’s rating can be. If $\sigma$ is initially set too high, the system will interpret the early game results as more variable or uncertain. This leads to larger rating adjustments, making the player’s rating fluctuate more. In contrast, setting $\sigma$ too low initially restricts the system from adjusting the rating freely. As a result, players, especially newcomers, may take much longer to reach a rating that accurately reflects their true skill level. This is particularly impactful in fast-changing or skill-evolving games, where a low initial variability setting could lead to underestimation of skill. We observe that Glicko’s prediction performance is consistent across both datasets, showing that its performance is less impacted by changes in its hyperparameters compared to Elo and Markov Chain. We theorize that this is due to the large number of games played for each model in both scenarios, allowing the system to adjust the rating deviation dynamically.
Elo shows high sensitivity to hyperparameter whereas Glicko demonstrates greater stability and accuracy which suggests that Glicko is a more reliable alternative for LLM evaluation.
5 Recommended Best Practices
Criteria Elo Markov Glicko Bradley-Terry
In sections 3 and 4, we outlined the desired properties of ranking systems and presented experiments assessing the performance of these systems in achieving these criteria. Based on these observations, this section puts these results into perspective and provides practical recommendations for selecting the most appropriate ranking algorithm. We outline which algorithms perform best under specific circumstances, offering guidance on their optimal use depending on the structure of the dataset, the matchup distribution, and the model evaluation goals. These recommendations, summarized in Table 4, aim to improve the reliability and applicability of classification systems in real-world LLM evaluations.
5.1 Challenges with Elo
An important result in this paper is that Elo, even with $>$ 1,000 permutations, struggles to achieve stable rankings A.5. This indicates that Elo’s dependency on permutations for ranking stabilization is inadequate, challenging previous studies Boubdir et al. (2023). Combined with the high sensitivity to hyperparameters and the average performance in preserving transitivity. We do not recommend ELO as an algorithm to rank LLM performance.
Best Practice 1. Elo’s high sensitivity to changes in its $K$ factor and instability (even with large permutations) make it unsuitable for efficient ranking of LLM performance in pairwise evaluation (Sections 4.4 and A.5).
5.2 Choose Bradley-Terry for Small Controlled Datasets
In section 4.3, we demonstrate that ranking algorithms exhibit varied performance based on battle conditions. When matches are evenly distributed, the ranking algorithms produce similar results. However, a heavily skewed distribution can negatively impact some methods such as Elo and Bradley-Terry. In this scenario, where matchups are balanced and the number of matches is small, objective findings show that Bradley-Terry is best due to its ability to preserve transitivity while also maintaining comparable prediction accuracy.
Best Practice 2. Bradley-Terry’s simplicity and lack of hyperparameters make it ideal for controlled evaluation settings, although its reliance on sufficient matchup data limits its effectiveness in unevenly distributed scenarios.
5.3 Choose Glicko for large uneven datasets
For larger datasets, especially those with uneven matchup distributions, the Bradley-Terry model encounters difficulties due to the rare event problem. This problem occurs when new models with high win rates, participate in fewer matchups than well-established models. As a result, these new models can be ranked disproportionately higher based on limited data, leading to biased rankings. In this scenario, the Glicko system provides a superior solution. Glicko utilizes a rating deviation parameter to account for uncertainty in a model’s ranking due to limited matchups. This allows the system to adjust the rankings more conservatively for models with fewer comparisons, preventing them from being ranked too high based only on a few matches. As data sets become larger and more uneven, Glicko’s ability to incorporate rating deviation ensures that the rankings accurately reflect the true performance throughout the dataset, even when new models appear with limited match histories. This flexibility makes Glicko particularly effective for large-scale LLM assessments where uneven matchup distributions are prevalent.
Best Practice 3. Glicko is effective for large data sets with uneven matches due to its rating deviation parameter, which manages uncertainty and prevents models with few matches from being overrated.
5.4 Bradley-Terry for other Dataset Types
For small, uneven datasets, Bradley-Terry’s simplicity and ability to model pairwise comparisons make it a robust choice, even when the data distribution is skewed. It does not require parameter estimation, which allows it to handle small datasets efficiently without the need for complex adjustments or hyperparameter tuning, making it ideal for situations where uneven distributions might otherwise require additional preprocessing.
In the case of large, evenly distributed datasets, Bradley-Terry is effective because of its scalability and ease of interpretation. For large data sets where the distribution of results is balanced, Bradley-Terry’s ability to produce clear, interpretable rankings without adding excessive computational overhead is a significant advantage. Although more complex models such as Glicko can offer finer granularity in large, uneven data cases, Bradley-Terry’s efficiency and interpretability make it a viable alternative when computational simplicity and transparency are prioritized.
Best Practice 4. Bradley-Terry’s scalability and interpretability make it a viable choice for large, evenly distributed datasets, offering clear rankings with minimal computational overhead.
6 Related Work
Pairwise ranking has recently emerged as a popular method for evaluating LLMs, as it allows for direct comparisons between models and provides a more nuanced assessment of their performance than traditional metrics Chiang et al. (2024). This evaluation involves presenting human judges with pairs of model outputs and asking them to select the better one. The outcomes of these pairwise comparisons are then aggregated into final rankings using ranking algorithms such as Elo Elo (1978). Despite their widespread use, there has been limited research on the applicability of these ranking algorithms to LLM evaluation. Works such as Boubdir et al. (2023) have critically analyzed the compatibility of Elo with LLM evaluations by examining how well it adheres to reliability and transitivity. However, Elo is not the only ranking algorithm used in LLM evaluation. There exists a number of other algorithms that have been proposed to rank in different contexts, such as the Bradley-Terry model Bradley and Terry (1952) and Glicko Glickman and Jones (1999), which have been shown to outperform Elo in certain scenarios, but their performance in the context of LLM evaluation has not been systematically evaluated. However, several factors affect the performance of these algorithms, such as the distribution of the data, the number of models that are compared, and the number of pairwise comparisons. As such, there is a need for a systematic comparison of ranking algorithms in the context of LLM evaluation to identify the most suitable algorithm for different scenarios which this paper aims to address.
7 Conclusion
8 Limitations
While our study comprehensively explores the ranking of Large Language Models (LLMs) through head-to-head combat, several limitations warrant discussion.
Scalability Constraints
One significant limitation arises from the scalability constraints inherent in pairwise evaluations. As the number of LLMs in the ecosystem expands, the number of required comparisons grows quadratically. This introduces computational challenges and resource constraints that may limit the feasibility of exhaustive head-to-head evaluations, particularly for larger sets of models.
Human Feedback Variability
Another concern is the variability in human feedback, on which our rankings are heavily based. Human judgments can be subjective and influenced by numerous factors, including the individual’s background, expertise, and contextual understanding. This variability can introduce noise into the ranking system, affecting the evaluations’ stability and reliability.
Appendix A Appendix
A.1 Glicko Algorithm Details